# Developer Perspectives on Licensing and Copyright Issues Arising from Generative AI for Software Development

### v3

@@ 3.2. Participant identification
We sought to learn from developers who used AI tools for code generation. To this end, and consistent with prior work (Joblin et al., 2017; Xia et al., 2023; Stalnaker et al., 2024; Liang et al., 2024; Li and Ahmed, 2023; Huang et al., 2021; Filippova and Cho, 2016; Jiang et al., 2017; Moraes et al., 2021; Constantino et al., 2023; He et al., 2023; Saroar and Nayebi, 2023; Liang et al., 2022), we used GitHub as a source to identify potential participants, following multiple steps.
We began by searching GitHub to identify repositories related to code generation via AI. We did this by developing a list of relevant tags, including “coding-assistant” and “ai-code-generator.” The full list of tags can be found in our replication package (ano, 2024). This resulted in 27 repositories, including open-source coding assistants such as “Tabby” (tabby, [n. d.]) and “cptX” (cptX, [n. d.]) as well as “full development” solutions such as “gpt-engineer” (gptengineer, [n. d.]) and “gpt-pilot” (gpt-pilot, [n. d.]). We included an additional three repositories we believed would be relevant: Meta’s CodeLlama, (B. Rozière et al., 2024), GitHub Copilot (GitHub, [n. d.]), and Codeium (codeium, [n. d.]). This brought our total repository count to 30. Next, using the GitHub API (GitHub REST API documentation, [n. d.]), we sought to identify developers who had shown interest in any of these repositories by forking, starring, watching, or contributing to them. We believe that these actions are highly correlated with familiarity with using GenAI tools for software development tasks, although we also acknowledge that our respondent sample may not be representative of all developers and may overrepresent those who frequently use GenAI tools (see Section 7). Lastly, to respect the privacy of developers, we filtered the collected information to eliminate users who did not provide publicly available contact information on their GitHub profile. This resulted in 30,659 public-facing email addresses, which we used to contact individuals seeking participation in our research study. All information reported in this paper, including demographic information, was contributed voluntarily by respondents who agreed to participate in response to our request.
@@ 3.3. Survey response collection and analysis
Survey responses to five open-ended questions were analyzed through a qualitative coding approach (Spencer, 2009). Two authors, both SE researchers, performed open-coding by independently assigning one or more codes to each response using a shared spreadsheet and codebook. Each annotator independently coded all 574 responses, adding new codes to the codebook as necessary. Once the initial coding was completed, the annotators met to settle disagreements and consolidate the set of codes. The disagreements encountered consisted of differing interpretations of the response, applying the same codes in different contexts, or using broader versus more specific codes. Through this process, all such disagreements were resolved, resulting in a final set of codes, definitions, and assignments for each response. Our replication package (ano, 2024) contains the final codes and definitions.
@@
We did not base our analysis on inter-rater agreements because multiple codes could be assigned to each response, and no list of codes existed before the start of coding (i.e., our approach was fundamentally inductive). However, we carefully followed best open-coding practices (Spencer, 2009), and we leveraged coders’ discussions to ensure the results’ reliability. All of the responses were coded in a single iteration by the two annotators. To mitigate agreement by chance and the introduction of bias, the labels assigned to each answer were reviewed by the annotators, including those without disagreement. The same annotators performed this review so as to be able to provide their understanding and rationale behind the labeling of each response. We avoided defining “umbrella” codes by creating complete definitions for each code and sharing them between the annotators during the review, as well as by assessing the appropriateness of codes during reconciliation of the annotators’ results such that codes that were too similar could be merged together and codes that were overbroad could be broken into multiple, more specific codes. This was done when the annotators observed multiple codes frequently being used to describe the same statement in a given answer or when a code was being applied to answers with significantly different meanings, respectively.
@@
Two open-ended questions solicited additional thoughts, comments, or explanations on respondents’ answers to a number of different questions. The disparity in responses prevented meaningful analysis with our previous coding method, so we applied a different strategy. One SE researcher read through all 574 responses for both questions and grouped them into categories and sub-categories based on their content by considering the respondent’s other answers as context. An additional researcher reviewed and validated these categorization decisions, discussing and resolving disagreements when needed. Lastly, as a separate task, a third author reviewed the quotes extracted for the paper, making sure none were miscategorized, misattributed, or taken out of context. For the closed-ended questions, we aggregated results using descriptive statistics and discussed them to resolve any disagreements. One author manually coded each response for questions that had an “Other” option (such as C2 and D4), so that this additional information could be effectively analyzed. Sections 4, 5, and 6 present the results of our analysis.
@@ 3.4. Participant demographics and background
Table 3. Top-10 country breakdown*
@@
Respondents self-selected the nature of their development work from a list that included proprietary software development (331), open source (247), and academia (117). The total number of responses to this question is greater than the number of complete and valid responses (574) because respondents could select more than one option. Respondents came from six continents, as indicated in Table 4, and 73 different countries, the top ten of which are shown in Table 3.
@@ 3.5. Follow-up Interviews
All survey participants were asked at the conclusion of the survey if they would be willing to be contacted for a follow-up interview. Those who consented by supplying their email address (382, 66.5%) made up our potential interview participant pool. Two authors then selected 22 potential interview candidates from this pool based on various criteria, including experience, likely legal jurisdiction based on location, and indicated views/perceptions on key legal issues, with the goal of interviewing participants representing a diversity of views. The full list of interview selection criteria can be found in our replication package (ano, 2024). Of the 22 candidates, seven responded to schedule an interview. The seven interviewees represented four continents (North America (3), Europe (2), South America (1), and Africa (1)) and a variety of backgrounds in software development, including working as a developer on proprietary software (6), contributing to OSS (4), and working in academia (1).
The duration of the interviews was between 20 and 30 minutes. The interviews were conducted using the Zoom video conferencing platform. Interview sessions were recorded and transcribed using OpenAI’s Whisper large-v3 model (Radford et al., 2023; Whisper, [n. d.]) to facilitate conversation analysis. A set of shared interview questions was iteratively derived through the collaboration of six SE researchers and one law researcher. This list of questions is available in our replication package (ano, 2024). Researchers were also free to ask follow-up or clarification questions during the interviews.
@@
Interview participants $RI_{31}$, $RI_{118}$, and $RI_{190}$ have experience in proprietary software development. $RI_{118}$ specifically works as a cyber security manager. $RI_{516}$ has contributed to open-source projects. $RI_{39}$, $RI_{149}$, and $RI_{431}$ work as closed-source software developers but have also contributed to open-source projects. $RI_{431}$ also has some experience writing software in an academic environment.
@@ 4. RQ1: Use of Generative AI technologies for software development
RQ1 aimed to discover how developers are using GenAI technologies for developing software. Although the developers surveyed had a range of perceptions about AI and copyright issues, most seemed to share the belief that the use of GenAI is now a regular part of everyday development activities.
In the following we present and discuss the findings for this RQ. We present the key conclusions that answer this research question as Findings 1-10. A summary of key findings can be found in our replication package (ano, 2024).
@@ 4.1.1. GenAI usage in software development
Almost all respondents (554) stated that they were generally familiar with the use of GenAI to produce source code. Of these, the vast majority—497 (90%)—indicated that they had in some capacity incorporated such tools into their development workflows. (Figure 3 displays the types of tasks mentioned by respondents.) The remaining 57 (10%) offered several reasons for not regularly using GenAI, including legal risk (8), personal preference (8), the lack of significant use cases warranting its use (6), reluctance to send sensitive information to a remote server (6), constraints imposed by an employer (6), and limited trust in the tooling (5).
Among those who had a personal preference against using GenAI, some developers conveyed through open-ended responses the sense that GenAI encourages monotony or lack of effort, despite the fact that GenAI is often used to automate repetitive and boring tasks, as previous work has indicated (Sergeyuk et al., 2024). $R_{188}$ stated, “I like to think for myself. [W]ithout [that], coding becomes boring and a lot of reading instead of reading and writing.” $R_{339}$ put it this way: “It’s better to write code on my own than to review code generated by AI.” The full list of reasons put forward by developers can be found in our replication package (ano, 2024).
Finding 1: While the vast majority of developers we surveyed (90%) were using GenAI tools, there were many reasons why some developers did not incorporate GenAI into their development workflows, including perceived legal issues, personal preference, information security concerns, and limitations placed by employers.
@@ 4.1.2. GenAI tools used by developers
Participants used various GenAI tools, including open-source, open-weight, and proprietary models. In total, developers reported using 64 distinct, named GenAI tools, the top 20 of which can be seen in Table 5. Developers were encouraged to identify all the tools they had used in their work. Of these, the top six (which were also provided in a list of choices, along with an “Other” option) were: GitHub Copilot (366), ChatGPT-4 (304), ChatGPT-3.5 (268), Bard (95), Claude (51), and Amazon CodeWhisperer (33). The next largest group (19) reported that they used open-source/open-weight models, with several such tools/models listed by multiple respondents, including Codellama (11), Mistral (9), Ollama (7), Llama2 (7), and Mixtral (6). Developers also used other proprietary models and services, including Codeium (16), Perplexity (12), Phind (10), and JetBrains AI (7), and six respondents indicated that they used custom-trained or in-house models for their development work. A full list of the tools developers reported using can be found in our replication package (ano, 2024).
Of the tools listed by respondents, 20 were designed to be integrated into an IDE. Many of these tools provide autocomplete suggestions for developers based on the context of surrounding code, but some also include Q/A chat functionality. Developers also reported using 18 distinct web-based tools, such as ChatGPT, Perplexity, and Hugging Face Chat (OpenAI, [n. d.]; Perplexity, [n. d.]; HuggingChat, [n. d.]), that require the user to supply more context than IDE-integrated tools. Respondents also indicated using four model infrastructures: Ollama, Oobabooga, llama.cpp, and Mozilla llamafile (Ollama, [n. d.]; Oobabooga, [n. d.]; Llama.cpp, [n. d.]; Hood, 2023). These infrastructures provide a streamlined means to run and train open-weight models locally. In total, respondents identified 15 distinct open-weight models/model families. (We note that the prevalence of open-weight models and open-source tools in this space has likely increased since our initial survey with the subsequent release of models such as Llama3 (Meta, 2024) and the open-sourcing of XAI’s Grok (XAI, 2024).) Three developers also noted the use of GenAI agents like gpt-engineer (gptengineer, [n. d.]) and metagpt (Hong et al., 2023), which allow models to self-prompt to accomplish more complex tasks. Lastly, four GenAI tools that were not designed for software development tasks, such as Figma (figma, [n. d.]) and GrammarlyGo (grammarly, [n. d.]), were mentioned. Figure 4 shows the usage of tools in the categories discussed here as reported by respondents.
@@
Finding 3: The ecosystem of GenAI tools used by developers was varied and extensive, comprising both open-source/open-weight and proprietary solutions. Developers primarily used web-based tools and those that could be integrated into their IDEs.
@@ 4.1.3. Benefits derived from GenAI usage
Respondents who had used GenAI as part of their workflow identified several benefits from its use. The greatest benefit mentioned by 288 developers (58%) was increased productivity, faster development, and efficiency. Other benefits included not having to spend time writing simple or boilerplate code (88), quick prototyping (42), debugging (37), faster, more accurate documentation (29), summarizing existing documentation as an alternative to reading it (28), providing inspiration (16), refactoring (14), use in writing test cases (5), domain logic support (4), and generating test data (4). A full list of the benefits identified by respondents can be found in our replication package (ano, 2024). We note that participants often mentioned tasks as benefits, likely indicating that GenAI improved the effectiveness, productivity, or ease of completion for those particular tasks.
Thirty respondents identified improvements in code quality and optimization over what human developers could achieve. $RI_{118}$ noted, “When you couple the lack of training [of the average developer] with generally one to three years’ experience, […] the fact is that GPT4 or many of these tools are actually better developers than they are. […] I would almost feel better if [a code comment] said I wrote this and then I got a code review from GPT4.” Similarly, developers related that GenAI could help them write software in unfamiliar languages (28) and libraries (27). During a follow-up interview, $RI_{190}$ described how they would “basically just writ[e] the code in Typescript and then just tell [ChatGPT] convert to Liquid.” Six respondents mentioned using GenAI specifically to port or translate code. $R_{90}$ summarized this benefit: “The syntax of various languages is [why it] can take hundreds of hours of use to establish [familiarity] in a new language, but AI significantly reduces the ramp-up time.”
Along the same lines, nineteen respondents noted that GenAI helped them to be better developers by enhancing their understanding of unfamiliar code. However, GenAI models are not limited to explaining source code or concepts. $R_{238}$ noted that they could also be used for “commit message generation [and ex]planation of changes in commit (summarization).”
@@ 4.1.4. Challenges and shortcomings encountered in using GenAI
Respondents also identified several challenges arising from using GenAI, the most significant of which was the generation of unhelpful, unwanted, or broken code (128). Additional challenges included hallucinations (64) (such as recommending non-existent libraries or method calls), problems arising from outdated training data (53), issues arising from tools not having access to the project’s wider context (52), which could be caused by insufficient context windows (40), and an inability to use GenAI to solve complex problems (43). $R_{148}$ elaborated: “AI can do simple code pretty well, it’s not that capable [of generating] more complex code, it’s incapable of generating well-architected code outside of extremely simple patterns.” $R_{33}$ provided an example: “In my experience, a simple Python/Django project that required some form customizations was just not possible for any AI tool at my disposal. The logical part, the thinking, and the experience gained from years of coding real-world applications and systems design are [not comparable to] an AI code generation tool like GPT or Gemini.” The full list of challenges reported by developers is included in our replication package (ano, 2024).
Twenty-five respondents noted that the ability to write good prompts is essential to effectively using GenAI tools. $R_{303}$ reported that “AI cannot complete the majority of tasks by itself, even simple ones[,] without time spent [by the developer on] prompt engineering.”
Writing effective prompts is not always an easy task, however. $R_{18}$ said, “It is hard to communicate what I want or what I’m trying to describe to the model.” If done correctly, however, prompting was also seen as a way to overcome other limitations of the technology. $R_{240}$ said, “There is a limit to the complexity that can be generated stand[ing] alone, but a human crafting intelligent prompts can still architect complex solutions using AI.”

### v5

@@ 3.2. Participant identification
We sought to learn from developers who used AI tools for code generation. To this end, and consistent with previous work (Joblin et al., 2017; Xia et al., 2023; Stalnaker et al., 2024; Liang et al., 2024; Li and Ahmed, 2023; Huang et al., 2021; Filippova and Cho, 2016; Jiang et al., 2017; Moraes et al., 2021; Constantino et al., 2023; He et al., 2023; Saroar and Nayebi, 2023; Liang et al., 2022), we used GitHub as a source to identify potential participants, following multiple steps.
We began by searching GitHub to identify repositories related to code generation via AI. We did this by developing a list of relevant tags, including “coding-assistant” and “ai-code-generator.” The full list of tags can be found in our replication package (ano, 2025). This resulted in 27 repositories, including open-source coding assistants such as “Tabby” (tabby, [n. d.]) and “cptX” (cptX, [n. d.]) as well as “full development” solutions such as “gpt-engineer” (gptengineer, [n. d.]) and “gpt-pilot” (gpt-pilot, [n. d.]). We included an additional three repositories we believed would be relevant: Meta’s CodeLlama, (B. Rozière et al., 2024), GitHub Copilot (GitHub, [n. d.]), and Codeium (codeium, [n. d.]). This brought our total repository count to 30. Next, using the GitHub API (GitHub REST API documentation, [n. d.]), we sought to identify developers who had shown interest in any of these repositories by forking, starring, watching, or contributing to them. We believe that these actions are highly correlated with familiarity with using GenAI tools for software development tasks, although we also acknowledge that our respondent sample may not be representative of all developers and may overrepresent those who frequently use GenAI tools (see Section 7). Lastly, to respect the privacy of developers, we filtered the collected information to eliminate users who did not provide publicly available contact information on their GitHub profile. This resulted in 30,659 public-facing email addresses, which we used to contact individuals seeking participation in our research study. All information reported in this paper, including demographic information, was contributed voluntarily by respondents who agreed to participate in response to our request.
@@ 3.3. Survey response collection and analysis
Survey responses to five open-ended questions were analyzed through a qualitative coding approach (Spencer, 2009). Two authors, both SE researchers, performed open-coding by independently assigning one or more codes to each response using a shared spreadsheet and codebook. Each annotator independently coded all 574 responses, adding new codes to the codebook as necessary. Once the initial coding was completed, the annotators met to settle disagreements and consolidate the set of codes. The disagreements encountered consisted of differing interpretations of the response, applying the same codes in different contexts, or using broader versus more specific codes. Through this process, all such disagreements were resolved, resulting in a final set of codes, definitions, and assignments for each response. Our replication package (ano, 2025) contains the final codes and definitions.
@@
We did not base our analysis on inter-rater agreements because multiple codes could be assigned to each response, and no list of codes existed before the start of coding (i.e., our approach was fundamentally inductive). However, we carefully followed best open-coding practices (Spencer, 2009), and we leveraged coders’ discussions to ensure the reliability of the results. All of the responses were coded in a single iteration by the two annotators. To mitigate agreement by chance and the introduction of bias, the labels assigned to each answer were reviewed by the annotators, including those without disagreement. The same annotators performed this review so as to be able to provide their understanding and rationale behind the labeling of each response. We avoided defining “umbrella” codes by creating complete definitions for each code and sharing them between the annotators during the review, as well as by assessing the appropriateness of codes during reconciliation of the annotators’ results such that codes that were too similar could be merged together and codes that were overbroad could be broken into multiple, more specific codes. This was done when the annotators observed multiple codes frequently being used to describe the same statement in a given answer or when a code was being applied to answers with significantly different meanings, respectively.
@@
Two open-ended questions solicited additional thoughts, comments, or explanations on respondents’ answers to a number of different questions. The disparity in responses prevented meaningful analysis with our previous coding method, so we applied a different strategy. One SE researcher read through all 574 responses for both questions and grouped them into categories and sub-categories based on their content by considering the respondent’s other answers as context. An additional researcher reviewed and validated these categorization decisions, discussing and resolving disagreements when needed. Lastly, as a separate task, a third author reviewed the quotes extracted for the paper, making sure none were miscategorized, misattributed, or taken out of context. For the closed-ended questions, we aggregated results using descriptive statistics and discussed them to resolve any disagreements. One author manually coded each response for questions that had an “Other” option (such as C3 and D4), so that this additional information could be effectively analyzed. Sections 4, 5, and 6 present the results of our analysis.
@@ 3.4. Participant demographics and background
Table 3. Top 10 country breakdown*
@@
Respondents self-selected the nature of their development work from a list that included proprietary software development (331), open source (247), and academia (117). The total number of responses to this question is greater than the number of complete and valid responses (574) because respondents could select more than one option. Respondents came from six continents, as indicated in Table 4, and 73 different countries (as self-reported), the top ten of which are shown in Table 3.
@@ 3.5. Follow-up Interviews
All survey participants were asked at the conclusion of the survey if they would be willing to be contacted for a follow-up interview. Those who consented by supplying their email address (382, 66.5%) made up our potential interview participant pool. Two authors then selected 22 potential interview candidates from this pool based on various criteria, including experience, likely legal jurisdiction based on location, and indicated views/perceptions on key legal issues, with the goal of interviewing participants representing a diversity of views. The full list of interview selection criteria can be found in our replication package (ano, 2025). Of the 22 candidates, seven responded to schedule an interview. The seven interviewees represented four continents (North America (3), Europe (2), South America (1), and Africa (1)) and a variety of backgrounds in software development, including working as a developer of proprietary software (6), contributing to OSS (4), and working in academia (1).
The duration of the interviews was between 20 and 30 minutes. The interviews were conducted using the Zoom video conferencing platform. Interview sessions were recorded and transcribed using OpenAI’s Whisper large-v3 model (Radford et al., 2023; Whisper, [n. d.]) to facilitate conversation analysis. A set of shared interview questions was iteratively derived through the collaboration of six SE researchers and one law researcher. This list of questions is available in our replication package (ano, 2025). Researchers were also free to ask follow-up or clarification questions during the interviews.
@@
Interview participants $RI_{31}$, $RI_{118}$, and $RI_{190}$ have experience in proprietary software development. ($RI_{118}$ currently works as a cybersecurity manager.) $RI_{516}$ has contributed to open-source projects. $RI_{39}$, $RI_{149}$, and $RI_{431}$ have experience in proprietary software development but have also contributed to open-source projects. $RI_{431}$ also has some experience writing software in an academic environment.
@@ 4. RQ1: Use of Generative AI technologies for software development
RQ1 aimed to discover how developers are using GenAI technologies to develop software. Although the developers surveyed had a range of perceptions about AI and copyright issues, most seemed to share the belief that the use of GenAI is now a regular part of everyday development activities. In the following, we present and discuss the findings for this RQ. We present the key conclusions that answer this research question as Findings 1-10. A summary of key findings can be found in our replication package (ano, 2025).
@@ 4.1.1. GenAI usage in software development
Almost all respondents (554) stated that they were generally familiar with the use of GenAI to produce source code. Of these, the vast majority—497 (90%)—indicated that they had incorporated such tools into their development workflows in some capacity. (Figure 3 displays the types of tasks mentioned by respondents.) The remaining 57 (10%) offered several reasons for not regularly using GenAI, including legal risk (8), personal preference (8), the lack of significant use cases that warrant its use (6), reluctance to send sensitive information to a remote server (6), constraints imposed by an employer (6), and limited trust in the tooling (5).
Among those who had a personal preference against using GenAI, some developers conveyed through open-ended responses the sense that GenAI encourages monotony or lack of effort, despite the fact that GenAI is often used to automate repetitive and boring tasks, as previous work has indicated (Sergeyuk et al., 2024). $R_{188}$ stated, “I like to think for myself. [W]ithout [that], coding becomes boring and a lot of reading instead of reading and writing.” $R_{339}$ put it this way: “It’s better to write code on my own than to review code generated by AI.” The full list of reasons put forward by developers can be found in our replication package (ano, 2025).
Finding 1: While the vast majority of developers we surveyed (90%) reported using GenAI tools, there were many reasons why some developers reported not incorporating GenAI into their development workflows, including perceived legal issues, personal preference, information security concerns, and limitations placed by employers.
@@ 4.1.2. GenAI tools used by developers
Participants used various GenAI tools, including open-source, open-weight, and proprietary models. In total, developers reported using 64 distinct GenAI tools, of which the top 20 can be seen in Table 5. Developers were encouraged to identify all the tools they had used in their work. Of these, the top six (which were also provided in a list of choices, along with an “Other” option) were: GitHub Copilot (366), ChatGPT-4 (304), ChatGPT-3.5 (268), Bard (95), Claude (51) and Amazon CodeWhisperer (33). The next largest group (19) reported that they used open-source/open-weight models, with several such tools/models listed by multiple respondents, including Codellama (11), Mistral (9), Ollama (7), Llama2 (7) and Mixtral (6). Developers also used other proprietary models and services, including Codeium (16), Perplexity (12), Phind (10), and JetBrains AI (7), and six respondents indicated that they used custom-trained or internal models for their development work. A full list of the tools that developers reported using can be found in our replication package (ano, 2025).
Of the tools listed by respondents, 20 were designed to integrate into an IDE. Many of these tools provide autocomplete suggestions for developers based on the context of surrounding code, but some also include Q/A chat functionality. Developers also reported using 18 distinct web-based tools, such as ChatGPT, Perplexity, and Hugging Face Chat (OpenAI, [n. d.]; Perplexity, [n. d.]; HuggingChat, [n. d.]), which require the user to provide more context than IDE-integrated tools. Respondents also indicated using four model infrastructures: Ollama, Oobabooga, llama.cpp, and Mozilla llamafile (Ollama, [n. d.]; Oobabooga, [n. d.]; Llama.cpp, [n. d.]; Hood, 2023). These infrastructures provide a streamlined means to run and train open-weight models locally. In total, respondents identified 15 distinct open-weight models/model families. (We note that the prevalence of open-weight models and open-source tools in this space has probably increased since our initial survey with the subsequent release of models such as Llama3 (Meta, 2024) and the open-source of XAI Grok (XAI, 2024).) Three developers also noted the use of GenAI agents like gpt-engineer (gptengineer, [n. d.]) and metagpt (Hong et al., 2023), which allow models to self-prompt to accomplish more complex tasks. Lastly, four GenAI tools that were not designed for software development tasks, such as Figma (figma, [n. d.]) and GrammarlyGo (grammarly, [n. d.]), were mentioned. Figure 4 shows the usage of tools in the categories discussed here as reported by respondents.
@@
Finding 3: The ecosystem of GenAI tools developers reported using was varied and extensive, comprising both open-source/open-weight and proprietary solutions. Developers primarily used web-based tools and those that could be integrated into their IDEs.
@@ 4.1.3. Benefits derived from GenAI usage
Respondents who had used GenAI as part of their workflow identified several benefits from its use. The greatest benefit mentioned by 288 developers (58%) was increased productivity, faster development, and efficiency. Other benefits included not having to spend time writing simple or boilerplate code (88), quick prototyping (42), debugging (37), faster, more accurate documentation (29), summarizing existing documentation as an alternative to reading it (28), providing inspiration (16), refactoring (14), use in writing test cases (5), domain logic support (4), and generating test data (4). A full list of the benefits identified by respondents can be found in our replication package (ano, 2025). We note that participants often mentioned tasks as benefits, likely indicating that GenAI improved the effectiveness, productivity, or ease of completion for those particular tasks.
Thirty respondents identified improvements in code quality and optimization over what human developers could achieve. $RI_{118}$ noted, “When you couple the lack of training [of the average developer] with generally one to three years’ experience, […] the fact is that GPT4 or many of these tools are actually better developers than they are. […] I would almost feel better if [a code comment] said I wrote this and then got a code review from GPT4.” Similarly, developers reported that GenAI could help them write software in unfamiliar languages (28) and libraries (27). During a follow-up interview, $RI_{190}$ described how they would “basically just writ[e] the code in Typescript and then just tell [ChatGPT] convert to Liquid.”
Six respondents mentioned using GenAI specifically to port or translate code. $R_{90}$ summarized this benefit: “The syntax of various languages is [why it] can take hundreds of hours of use to establish [familiarity] in a new language, but AI significantly reduces the ramp-up time.” Along the same lines, nineteen respondents noted that GenAI helped them become better developers by improving their understanding of unfamiliar code. However, GenAI models are not limited to explaining source code or concepts. $R_{238}$ noted that they could also be used for “commit message generation [and ex]planation of changes in commit (summarization).”
@@ 4.1.4. Challenges and shortcomings encountered in using GenAI
Respondents also identified several challenges that arise from the use of GenAI, the most significant being the generation of unhelpful, unwanted, or broken code (128). Additional challenges included hallucinations (64) (such as recommending nonexistent libraries or method calls), problems arising from outdated training data (53), issues arising from tools not having access to the project’s wider context (52), which could be caused by insufficient context windows (40), and an inability to use GenAI to solve complex problems (43). $R_{148}$ elaborated: “AI can do simple code pretty well, it’s not that capable [of generating] more complex code, it’s incapable of generating well-architected code outside of extremely simple patterns.” $R_{33}$ provided an example: “In my experience, a simple Python/Django project that required some form customizations was just not possible for any AI tool at my disposal. The logical part, the thinking, and the experience gained from years of coding real-world applications and systems design are [not comparable to] an AI code generation tool like GPT or Gemini.” The full list of challenges reported by developers is included in our replication package (ano, 2025).
Twenty-five respondents noted that the ability to write good prompts is essential to effectively using GenAI tools. $R_{303}$ reported that “AI cannot complete the majority of tasks by itself, even simple ones[,] without time spent [by the developer on] prompt engineering.” Writing effective prompts is not always an easy task, however. $R_{18}$ said, “It is hard to communicate what I want or what I’m trying to describe to the model.” If done correctly, however, prompting was also seen as a way to overcome other limitations of the technology. $R_{240}$ said, “There is a limit to the complexity that can be generated stand[ing] alone, but a human crafting intelligent prompts can still architect complex solutions using AI.”